Question goes here. Write clearly!How do you choose your x and y values to calculte for your orders pair?
Answer by nyc_function(2741)

SAMPLE:
Find two points in the (x,y) for y = x + 3.
Select different values for x. Any value at all.
I will let x = 0 and then let x = 2.
When x = 0, y = 0 + 3 becomes 3 making y = 3.
Our first point is (0,3).
When x = 2, y = 2 + 3 becomes 5 making y = 5.
Our second point is (2,5).
Understand?